# Sampling Strategies in Bayesian Inversion: A Study of RTO and Langevin Methods

### v1

@@ Abstract
This paper studies two classes of sampling methods for the solution of inverse problems, namely Randomize-Then-Optimize (RTO), which is rooted in sensitivity analysis, and Langevin methods, which are rooted in the Bayesian framework. The two classes of methods correspond to different assumptions and yield samples from different target distributions. We highlight the main conceptual and theoretical differences between the two approaches and compare them from a practical point of view by tackling two classical inverse problems in imaging: deblurring and inpainting. We show that the choice of the sampling method has a significant impact on the quality of the reconstruction and that the RTO method is more robust to the choice of the parameters.
@@ 1 Introduction
where $x^{*}$ denotes the reconstruction. The objective function consists of a data-fidelity term $f(Ax,y)$ and a regularization term $g(x)$. The data-fidelity term measures the discrepancy between the observation $y$ and the the model output $Ax$ whereas the regularization term $g(x)$ penalizes images with undesirable properties, e.g., images outside a neighborhood of some sub-manifold. In many cases, the minimization problem in (2) can be solved efficiently using advanced optimization methods. However, a solution to the problem in (2) does not provide information about the inherent uncertainty that may arise because of noisy measurements, discretization, and/or model errors.
@@
The purpose of this paper is to contrast and compare two classes of sampling methods for solving inverse problems within the Bayesian framework, namely Randomize-Then-Optimize (RTO) [4] and Langevin sampling methods [34, 13, 14]. On one hand, RTO is deeply rooted within the variational framework as it solves a perturbed optimization problem in order to generate one sample. On the other hand, Langevin methods stem from the discretization of a stochastic differential equation (SDE) whose solution is the target distribution. The two classes of methods are similar in many ways, but they are based on different principles and assumptions, and hence they lead to different densities. This is illustrated in Figure 1, which shows different solution densities obtained from an observation $y=x+n$ where $n$ is standard normal and $x\in[a,b]$ is an unknown parameter. Adopting a uniform prior on $[a,b]$ leads to a truncated Gaussian posterior. The Langevin approach requires smoothness and leads to a smooth approximate truncated Gaussian posterior. In contrast, the RTO approach yields a density that is a mixture of a truncated Gaussian distribution and a distribution on the boundary of the interval. The example highlights some key differences: unlike the smooth approximate truncated Gaussian posterior, the RTO density assigns non-zero probability to the boundary of the interval and has compact support. This RTO density is associated with an implicit prior that is supported on $[a,b]$ and depends on the observation $y$. Consequently, it violates the typical Bayesian assumption that the prior is independent of the observation. Using somewhat unconventional terminology, we will refer to the RTO density as the RTO posterior.
@@ 2.1 Randomize-Then-Optimize (RTO) sampling method
$x^{\dagger}\in\operatorname*{argmin}_{x}\,\left\{\frac{1}{2\sigma^{2}}\|\mathcal{A}(x)-\hat{y}\|_{2}^{2}+g(x)\right\},\qquad g(x)=(x-x_{0})^{T}S^{-1}(x-x_{0}),$ (4)
where the perturbed data $\hat{y}$ is drawn from the Gaussian pdf $\mathcal{N}(y,\sigma^{2}I)$. For a linear inverse problem, this procedure yields independent samples from the posterior and reduces to a technique for Gaussian sampling proposed in [27]. When dealing with a nonlinear forward operator $\mathcal{A}$, the problem (4) transforms into a nonlinear least-squares problem, and RTO generates samples from an approximate posterior. However, samples from the exact posterior can be obtained by incorporating a Metropolis-Hastings step.
@@ 2.3 Conceptual differences between RTO and MYULA
Comparing the computational costs, we need to solve an optimization problem, i.e., computing (6b), in order to generate an RTO sample. However, with MYULA we obtain one sample at each iteration that involves the computation of a proximal operator, i.e., computing (10a). In RTO, as the perturbations are independent, samples are independent. Furthermore, we can parallelize RTO to generate several samples at the same time. MYULA, as all Markov chain Monte Carlo (MCMC) methods, require a burn-in period in order for the Markov chain to enter its stationary regime and actually produce samples from its stationary density. The length of the burn-in period is unknown a-priori. In addition, samples generated from MYULA as well as other Langevin methods suffer from correlations as they are generated by using the previous sample. Further, this type of methods cannot be parallelized, contrary to RTO.
Finally, we will compare the flexibility on the extension to model parameter selection. Often the posteriors include some unknown parameters, e.g. $\sigma$ in (4), $\gamma$ in (5) and $\alpha$ in (9a). The choice of those parameters can have great impact on the posterior. RTO allows us to perform automatic parameter selection by considering a hierarchical model, i.e. we add prior distributions on the hyperparameters $\lambda=1/\sigma^{2}$ and $\gamma$ and consider the posterior
@@
with $g(x)$ given in (5).
Algorithm 1 summarizes the procedure to sample from such a hierarchical model within the RTO framework.
@@
Selecting posterior parameters in the Langevin-based methods is much more difficult. As far as we know, no Langevin-based methods can update or simultaneously estimate the model parameters. Recently, Vidal et al. proposed to perform parameter selection based on maximum likelihood estimation [37, 12], which can simultaneously estimate multiple model parameters and is theoretically well-founded. But the procedure has to be implemented offline, i.e., the parameters have to be predetermined before sampling.
@@ 3 Experimental study
We generate 1000 RTO samples. To solve the minimization problem (4) with $g(x)$ defined in (13) in RTO, we apply the alternating direction method of multipliers (ADMM) with the stopping criterion suggested in [7, Eq (3.12)]. According to the theories in [15, 16], to guarantee a well-defined posterior density we need to add a term $\frac{\alpha}{2}\|x\|_{2}^{2}$ for the rank-deficient $A$, which is the case here, and we set $\alpha=10^{-8}$ to ensure little impact on the results.
@@
where $[\gamma\|\nabla.\|_{1,1}]_{\alpha_{1}}$ and $[i_{c}]_{\alpha_{2}}$ denote the Moreau-Yoshida envelopes of the two terms in $g$, respectively. To apply MYULA defined in (9a), we need $\operatorname{prox}_{\gamma\|\nabla.\|_{1,1}}^{\alpha_{1}}$. However, since $\operatorname{prox}_{\gamma\|\nabla.\|_{1,1}}^{\alpha_{1}}$ has no closed-form expression, we apply the first-order primal-dual algorithm introduced in [10] to estimate it. In the following, we set the number of iterations to estimate this proximal operator to $n_{pd}=50$. We initialize MYULA at the data $y$ and run $10^{6}$ iterations with the first $2.5\times 10^{4}$ iterations as burn-in. In addition, we set the thinning parameter to $250$, i.e., we store a sample every $250$ iterations. It ends up with $3900$ samples in total. As suggested in [14], we set $\alpha_{1}=\alpha_{2}=1/(\|A^{T}A\|/\sigma^{2})$ and $\delta=1/(2/\alpha_{1}+2/\alpha_{2}+2\|A^{T}A\|/\sigma^{2})$.
@@ 3.1 Deblurring
PSNR=25.69/SSIM=0.79 PSNR=24.97/SSIM=0.72 PSNR=25.17/SSIM=0.77
@@
For the computational time, it took around $6$ hours to generate 1000 RTO samples and around $35$ hours to generate 3900 MYULA samples after burn-in and thinning on an Intel(R) Core(TM) i9-9980HK CPU @ 2.40GHz. To reduce the computational cost, we can decrease the accuracy when solving the optimization problems in the sampling methods, i.e., the computation of $\operatorname{prox}_{\gamma\|\nabla.\|_{1,1}}^{\alpha_{1}}$ in MYULA and MAP on the perturbed data in RTO. In Figure 9 we show the results produced by MYULA when setting $n_{pd}=10$ for Simpson. In this case, the running time is reduced to around $10$ hours. Comparing with the results in Figure 4 and Figure 5 where we set $n_{pd}=50$, we can see that MMSE is more noisy and the uncertainty is not well discovered. In addition, the ACFs figure shows that we need many more iterations in order to obtain uncorrelated samples. In RTO we change tol in ADMM from $10^{-4}$ to $10^{-2}$ and show the results in Figure 10. It is clear that MMSE is nearly identical to the one in Figure 4. However, the standard deviation close to the edges are slightly smaller than the ones in Figure 5. It means that reasonably reducing the solution accuracy has a minor influence on the MMSE estimate, but can lead to insufficiently explored uncertainty.
@@ 3.2 Inpainting
PSNR=35.01/SSIM=0.94 PSNR=33.86/SSIM=0.89 PSNR=34.99/SSIM=0.95
@@
PSNR=30.77/SSIM=0.92 PSNR=30.43/SSIM=0.89 PSNR=30.63/SSIM=0.92
@@
Finally, we consider the efficiency of both methods. The RTO experiments took around $10$ hours while for MYULA they took around $35$ hours.
In Figure 14 we show the evolution of the ACFs for both methods. We computed them in the pixel domain, where the forward operator $A$ is diagonal, as we assume that the likelihood imposes the posterior covariance shape. It is obvious that samples generated by RTO are independent, but samples generated with MYULA exhibit correlations. For Traffic, the complete Markov chain associated with MYULA needs more than $6000$ iterations in order to eventually achieve uncorrelated samples.
@@ 3.3 Automatic parameter selection
In Figure 15 we show the results from the same deblurring problem as in Section 3.1 but with slightly different $g(x)$ comparing with (13). Here we use the non-negativity constraint instead of the box constraint, i.e., $g(x)=\gamma\|\nabla x\|_{1,1}+i_{\mathbb{R}^{+}}(x)$, since in this case the conditional distribution $𝛾\pi_{\bbgamma|\mathbbm{x},\mathbbm{y}}(\gamma)$ can be efficiently sampled. Comparing the results shown in Figure 15 and in Figures 4 and 5, we can see that they are very similar visually as well as quantitatively. In Table 2, we list the means and the standard deviations of $\lambda$ and $\gamma$ obtained from the hierarchical Gibbs sampler. It is clear that the estimates for both $\gamma$ and $\lambda$ are comparable with the ones used in Section 3.1. In the end, Figure 16 shows the ACFs of the image samples and trace plots of $\lambda$ and $\gamma$. Due to the Gibbs sampler, we cannot expect independent samples, but we notice that ACFs still decay to zero immediately. Further, the trace plots show very good mixing in samples.
@@ 4 Conclusion
RTO is derived from the sensitivity analysis framework and samples from a target distribution by solving perturbed optimization problems where the perturbation occurs in the data space. The RTO target density can have non-zero probability mass on subsets of measure zero. However, it is not anchored in the Bayesian framework as the target density corresponds to an implicit prior that depends on the observed data. In addition, RTO can be incorporated into a hierarchical model in order to perform automatic parameter selection. The main limitation of RTO is that it has only been characterized for the posteriors with Gaussian likelihoods and polyhedral hypograph log-prior. In contrast, MYULA is firmly rooted in the Bayesian framework and is applicable to a broader range of posteriors. Although it samples the posterior density approximately, the distribution behind samples can be characterized with respect to the posterior density. Similar as other Langevin methods, MYULA suffers from typical MCMC drawbacks, such as slow convergence, correlated samples, etc. Through two classical imaging inverse problems: deblurring and inpainting, we compared RTO and MYULA numerically with particular attention to computational cost. Both methods produced accurate results for deblurring, but MYULA struggled with severely ill-posed problems like inpainting. Additionally, while RTO concentrates the sample mass around the MMSE estimate, MYULA results in a more dispersed distribution.

### v3

@@ Abstract
This paper studies two classes of sampling methods for the solution of inverse problems, namely Randomize-Then-Optimize (RTO), which is rooted in sensitivity analysis, and Langevin methods, which are rooted in the Bayesian framework. The two classes of methods correspond to different assumptions and yield samples from different target distributions. We highlight the main conceptual and theoretical differences between the two approaches and compare them from a practical point of view by tackling two classical inverse problems in imaging: deblurring and inpainting. We show that the choice of the sampling method has a significant impact on the reconstruction and the proposed uncertainty.
@@ 1 Introduction
where $x^{*}$ denotes the reconstruction. The objective function consists of a data-fidelity term $f(Ax,y)$ and a regularization term $g(x)$. The data-fidelity term measures the discrepancy between the observation $y$ and the model output $Ax$ whereas the regularization term $g(x)$ penalizes images with undesirable properties, e.g., images outside a neighborhood of some sub-manifold. In many cases, the minimization problem in (2) can be solved efficiently using advanced optimization methods. However, a solution to the problem in (2) does not provide information about the inherent uncertainty that may arise because of noisy measurements, discretization, and/or model errors.
@@
The purpose of this paper is to contrast and compare two classes of sampling methods for solving inverse problems within the Bayesian framework, namely Randomize-Then-Optimize (RTO) [2] and Langevin sampling methods [3, 4, 5]. On one hand, RTO is deeply rooted within the variational framework as it solves a perturbed optimization problem in order to generate one sample. On the other hand, Langevin methods stem from the discretization of a stochastic differential equation (SDE) whose blue stationary distribution is the distribution of interest. The two classes of methods are similar in many ways, but they are based on different principles and assumptions, and hence they lead to different densities. This is illustrated in Figure 1, which shows different solution densities obtained from an observation $y=x+n$ where $n$ is standard normal and $x\in[a,b]$ is an unknown parameter. Adopting a uniform prior on $[a,b]$ leads to a truncated Gaussian posterior. The Langevin approach requires smoothness and leads to a smooth approximate truncated Gaussian posterior. In contrast, the RTO approach yields a density that is a mixture of a truncated Gaussian distribution and a distribution on the boundary of the interval. The example highlights some key differences: unlike the smooth approximate truncated Gaussian posterior, the RTO density assigns non-zero probability to the boundary of the interval and has compact support. This RTO density is associated with an implicit prior that is supported on $[a,b]$ and depends on the observation $y$. Consequently, it violates the typical Bayesian assumption that the prior is independent of the observation. Using somewhat unconventional terminology, we will refer to the RTO density as the RTO posterior.
@@ 2.1 Randomize-Then-Optimize (RTO) sampling method
$x^{\dagger}\in\operatorname*{argmin}_{x}\,\left\{\frac{1}{2\sigma^{2}}\|Q^{T}[\mathcal{A}(x)-\hat{y}]\|_{2}^{2}+g(x)\right\},\qquad g(x)=(x-x_{0})^{T}S^{-1}(x-x_{0}),$ (4)
where the perturbed data $\hat{y}$ is drawn from the Gaussian pdf $\mathcal{N}(y,\sigma^{2}I)$ and $Q$ is a matrix determining the tangent space of the posterior mode. It results from a $QR$ decomposition of the Jacobian of $\mathcal{A}$, $J_{\mathcal{A}}$, evaluated at $x^{*}=\operatorname*{argmin}_{u}\|Au-y\|_{2}^{2}/2\sigma^{2}+g(u)$. The idea of RTO in the non-linear case is to generate samples from a Gaussian distribution and to project them onto a manifold in a direction orthogonal to the tangent space of $J_{\mathcal{A}}(x^{*})$. This process generates samples from an approximate posterior with computable density $\hat{\pi}(x|y)$. Then samples from the exact posterior can be obtained by adding a Metropolis-Hastings step [2, 26]. For a linear inverse problem, where $\mathcal{A}(x)=Ax$ and $Q=I$, this procedure yields independent samples from the posterior and reduces to a technique for Gaussian sampling proposed in [27].
@@ 2.3 Conceptual differences between RTO and MYULA
Comparing the computational costs, we need to solve an optimization problem, i.e., computing (6b), in order to generate an RTO sample. However, with MYULA we obtain one sample at each iteration that involves the computation of a proximal operator, i.e., computing (10a). In RTO, as the perturbations are independent, samples are independent. Furthermore, we can parallelize RTO to generate several samples at the same time. MYULA, as all Markov chain Monte Carlo (MCMC) methods, require a burn-in period in order for the Markov chain to enter its stationary regime and actually produce samples from its stationary density. The length of the burn-in period is unknown a priori. In addition, samples generated from MYULA as well as other Langevin methods are correlated as they are generated by using the previous sample. Although expected, it can have a negative impact on the accuracy of the Monte Carlo estimates. Furthermore, to address the common convergence speed issues of MCMC methods and accelerate their convergence, Jacob et al. proposed in [35] to compute Monte Carlo estimates by coupling Markov chains running in parallel. However, their practical implementations in a very high-dimensional setting remains a challenge.
Finally, we will compare the flexibility on the extension to model parameter selection. Often the posteriors include some unknown parameters, e.g. $\sigma$ in (4), $\gamma$ in (5) and $\alpha$ in (9a). The choice of those parameters can have a great impact on the posterior [36, 37]. RTO allows us to perform automatic parameter selection by considering a hierarchical model, i.e. we add prior distributions on the hyperparameters $\lambda=1/\sigma^{2}$ and $\gamma$ and consider the posterior
@@
with $g(x)$ given in (5). Algorithm 1 summarizes the procedure to sample from such a hierarchical model within the RTO framework.
@@
[…]
@@ 3 Experimental study
We generate 1000 RTO samples. To solve the minimization problem (4) with $g(x)$ defined in (13) in RTO, we apply the alternating direction method of multipliers (ADMM) with the stopping criterion suggested in [40, Eq (3.12)]. According to the theories in [32, 31], to guarantee a well-defined posterior density we need to add a term $\frac{\alpha}{2}\|x\|_{2}^{2}$ for the rank-deficient $A$, which is the case here, and we set $\alpha=10^{-8}$ to ensure little impact on the results.
@@
where $[\gamma\|\nabla.\|_{1,1}]_{\alpha_{1}}$ and $[i_{c}]_{\alpha_{2}}$ denote the Moreau-Yoshida envelopes of the two terms in $g$, respectively. To apply MYULA defined in (9a), we need $\operatorname{prox}_{\gamma\|\nabla.\|_{1,1}}^{\alpha_{1}}$. However, since $\operatorname{prox}_{\gamma\|\nabla.\|_{1,1}}^{\alpha_{1}}$ has no closed-form expression, we apply the first-order primal-dual algorithm introduced in [41] to estimate it. In the following, we set the number of iterations to estimate this proximal operator to $n_{pd}=50$. We initialize MYULA at the data $y$ and run MYULA for $2.5\times 10^{5}$ iterations after $2.5\times 10^{4}$ burn-in iterations. As suggested in [5], we set $\alpha_{1}=\alpha_{2}=1/(\|A^{T}A\|/\sigma^{2})$ and $\delta=1/(2/\alpha_{1}+2/\alpha_{2}+2\|A^{T}A\|/\sigma^{2})$.
@@ 3.1 Deblurring
PSNR=25.69/SSIM=0.79 PSNR=24.98/SSIM=0.72 PSNR=25.17/SSIM=0.77
@@
For the computational time, it took around $6$ hours to generate 1000 RTO samples and around $10$ hours to generate $2.5\times 10^{5}$ MYULA samples after burn-in on an Intel(R) Core(TM) i9-9980HK CPU @ 2.40GHz. The cost of generating a sample using RTO is higher than with MYULA. To reduce the computational cost, we can decrease the accuracy when solving the optimization problems in the sampling methods, i.e., the computation of $\operatorname{prox}_{\gamma\|\nabla.\|_{1,1}}^{\alpha_{1}}$ in MYULA and MAP on the perturbed data in RTO. In Figure 9 we show the results produced by MYULA when setting $n_{pd}=10$ for Simpson. In this case, the running time is reduced to around $3$ hours. Comparing with the results in Figure 4 and Figure 5 where we set $n_{pd}=50$, we can see that MMSE is more noisy and the uncertainty is not well discovered. In addition, the ACFs figure shows that we need many more iterations in order to obtain uncorrelated samples. In RTO we change tol in ADMM from $10^{-4}$ to $10^{-2}$ and show the results in Figure 10. It is clear that MMSE is nearly identical to the one in Figure 4. However, the standard deviation close to the edges are slightly smaller than the ones in Figure 5. It means that reasonably reducing the solution accuracy has a minor influence on the MMSE estimate, but can lead to insufficiently explored uncertainty.
@@ 3.2 Inpainting
PSNR=35.01/SSIM=0.94 PSNR=33.85/SSIM=0.89 PSNR=34.99/SSIM=0.95
@@
PSNR=30.77/SSIM=0.92 PSNR=30.44/SSIM=0.89 PSNR=30.63/SSIM=0.92
@@
Finally, we consider the efficiency of both methods. The experiments took around $10$ hours both RTO and MYULA.
In Figure 14 we show the evolution of the ACFs for both methods. We computed them in the pixel domain, where the forward operator $A$ is diagonal, as we assume that the likelihood imposes the posterior covariance shape. It is obvious that samples generated by RTO are independent, but samples generated with MYULA exhibit correlations. For Traffic, the complete Markov chain associated with MYULA needs more than $6000$ iterations in order to eventually achieve uncorrelated samples. However, recall that generating one sample with RTO is approximately 250 times more expensive than with MYULA.
@@ 3.3 Automatic parameter selection
In Figure 15 we show the results from the same deblurring problem as in Section 3.1 but with slightly different $g(x)$ comparing with (13). Here we use the non-negativity constraint instead of the box constraint, i.e., $g(x)=\gamma\|\nabla x\|_{1,1}+i_{\mathbb{R}^{+}}(x)$, since in this case the conditional distribution $𝛾\pi_{\bbgamma|\mathbbm{x},\mathbbm{y}}(\gamma)$ can be efficiently sampled. Comparing the results shown in Figure 15 and in Figures 4 and 5, we can see that they are very similar visually as well as quantitatively. In Table 2, we list the means and the standard deviations of $\lambda$ and $\gamma$ obtained from the hierarchical Gibbs sampler. It is clear that the estimates for both $\gamma$ and $\lambda$ are comparable with the ones used in Section 3.1. In the end, Figure 16 shows the ACFs of the image samples and trace plots of $\lambda$ and $\gamma$. Due to the Gibbs sampler, we cannot expect independent samples, but the trace plots show very good mixing properties. It indicates that the samples produced by our augmented model are practically uncorrelated [26, 45].
@@ 4 Conclusion
RTO is derived from the sensitivity analysis framework and samples from a target distribution by solving perturbed optimization problems where the perturbation occurs in the data space. The RTO target density can have non-zero probability mass on subsets of measure zero. However, it is not anchored in the Bayesian framework as the target density corresponds to an implicit prior that depends on the observed data. In addition, RTO can be incorporated into a hierarchical model in order to perform automatic parameter selection. The main limitation of RTO is that it has only been characterized for the posteriors with Gaussian likelihoods and polyhedral hypograph log-prior. Its efficiency also depends on the ability to efficiently solve perturbed MAP optimization problems. In contrast, MYULA is firmly rooted in the Bayesian framework and is applicable to a broader range of posteriors. Generating a single sample with MYULA is rather cheap. Although it samples the posterior density approximately, the distribution behind samples can be characterized with respect to the posterior density. Through two classical imaging inverse problems: deblurring and inpainting, we compared RTO and MYULA numerically with particular attention to computational cost. Both methods produced accurate results for deblurring, but MYULA struggled with severely ill-posed problems like inpainting. Additionally, while RTO concentrates the sample mass around the MMSE estimate, MYULA results in a more dispersed distribution.